OpenStudy (anonymous):

John began his job making $20 the first day. After that he was paid $6.00 per hour. Y = 6x + 20. What is the intercept in the problem? Help is appreciated :)

OpenStudy (anonymous):

20 would be the intercept

OpenStudy (anonymous):

the 20 dollars he made the first day

OpenStudy (anonymous):

6 would be the slope, correct?

OpenStudy (anonymous):

yup

OpenStudy (mikurout):

Use general form of equation y = mx +c

OpenStudy (anonymous):

6 is the slope man

OpenStudy (camerondoherty):

The form is like this: Y=Mx+B m = slope b = y intercept

OpenStudy (anonymous):

B would be the intercept

OpenStudy (camerondoherty):

since b is your y intercept then 20 is your y intercept

OpenStudy (anonymous):

yup but u have to explain it so the person doesn't panic about the difficulty of the problem

OpenStudy (acxbox22):

everyone up here is right 20 is your y intercept as it is in the form y=mx+b

OpenStudy (anonymous):

Thank you everybody. Wish I could give you all medals. I appreciate it. I'm not the sharpest tool in the shed.

OpenStudy (anonymous):

WHO GOT THE MEDAL?

OpenStudy (acxbox22):

well you are not the bluntest either ;)

OpenStudy (acxbox22):

you know you can close the question right?